The hypotenuse of a right triangle measures 10m. One leg of the triangle is 2m longer then the other. Find the lengths of the legs.

How would you solve this equation?
Answer #1
First of all, you have to write the equation. Call the shorter leg of the triangle x.


x^2 + (x+2)^2 = 10^2 = 100

2x^2 + 4x + 4 = 100
x^2 + 2x -48 = 0
(x+8)(x-6) = 0

Solve for x etc
answered by: dashawn
Answer #2
thank you :)
answered by: Popularity